OpenStudy (anonymous):

module two quiz anyone???Assessment 02.04

OpenStudy (anonymous):

Assessment 02.04 Module Two Quiz 1.Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. 2.Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x 3.Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10, and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem. 4.A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

TheSmartOne (thesmartone):

Do you have aspecific question?

OpenStudy (anonymous):

she gave four @TheSmartOne

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

@cadiecatexo do u want a complete solution?

OpenStudy (anonymous):

yea yes please

OpenStudy (anonymous):

if u guys can help me with one or two questios id so appreciate it!!

OpenStudy (anonymous):

1. profit on 1 flower= $(40/100) * 0.50 profit on n flowers= $(40/100) * 0.50 * n $500 = $(40/100) * 0.50 * n n = 2500 flowers

OpenStudy (anonymous):

got it? @cadiecatexo

OpenStudy (anonymous):

thanks so much!!

OpenStudy (anonymous):

what question

OpenStudy (anonymous):

question 1 profit=revenue - cost p = 0.20x 500 = 0.20x then p =500 now you need to solve for x solve x = 2,500 angela will have to buy 2500 flower to get the profit of $500

OpenStudy (anonymous):

question 2 max has an error in steps 3 also 4*-1 doesnt equal -1. step 3 would be, 9x + 18 = 28 x- 4 not 9x + 18 - 28 x -1 the correct solution is 3(3x+6)=4(7x-1) 9x +18 = 28x-4 18=19x-4 22/19=x check the answer 3(22/19+6)/8 7(22/19-1)/6 45/38

OpenStudy (anonymous):

question 3 10A+6(5)=80 10A +$30=$80 10A=$80-$30 10A=$50 A=$50/$10 A=$5

OpenStudy (anonymous):

i got 44/50 i got the 4 one wrong